# Quantum walks, quantum cellular automata, quantum field theory, symmetry, interactions

### v1

@@ 6.1 Translation Symmetry
Translation symmetry implies that the evolution operator is invariant under shifts of the spatial lattice. In a 3D cubic lattice of size $N$ with periodic boundaries, we can define shift operators $S_{X,Y,Z}$ that move the states of the local subsystems to the site $\Delta x$ away in the positive $X$, $Y$ or $Z$ direction, respectively, wrapping around at the periodic boundaries. $S^{\dagger}_{X,Y,Z}$ obviously represents a shift in the negative direction. Translation symmetry was extensively studied in the early QCA literature Watrous95 ; Meyer96a ; Meyer96 . On the cubic lattice these shifts all commute.
If $S$ is a product of any number of these unit shifts $S_{X,Y,Z}$ or $S^{\dagger}_{X,Y,Z}$ then translation symmetry of the QCA evolution unitary $\hat{U}$ implies
@@
Since $\hat{U}$ and $S$ commute, we can find a basis of eigenstates of $\hat{U}$ that are eigenstates $S$. This is what allows the momentum picture that we used both for QWs and QCAs in this paper.
@@ 6.3 Parity
$\mathcal{P}Q\mathcal{P}=Q,\ \ \mathcal{P}P^{\pm}_{j}\mathcal{P}=P^{\mp}.$ (78)
The simplest way of satisfying these conditions is if $\mathcal{P}$ acts on the internal space like $Q$, and if $Q$ anticommutes with the operators $\Delta P_{j}=P_{j}^{+}-P_{j}^{-}$. This anticommutation plays an important role in the emergence of the Dirac equation in the long-wavelength limit of QWs MlodinowBrun18 ; MlodinowBrun20 ; BrunMlodinow20 ; MlodinowBrun21 .
@@ 6.4 Time-Reversal Symmetry and Negative Energies
Time-reversal symmetry captures the idea that time evolution looks the same forward and backward. Classically, this would be illustrated by reversing the momenta of all particles while leaving their positions unchanged; this is like “running the movie backwards.” In quantum mechanics we also need to flip the spins of particles. This symmetry is very important to the tension between locality of interactions, one the one hand, and positive energy on the other, as we shall see in Section 7.
@@
where $K$ is complex conjugation ($K{|\psi\rangle}={|\psi\rangle}^{*}$) and $V$ is a unitary operator. Plugging this form into Eq. (79) we get
@@
Suppose that the evolution unitary $\hat{U}$ has eigenstates $\{{|\phi_{m}\rangle}\}$ with eigenvalues $\{e^{i\phi_{m}}\}$. Then $\hat{U}^{*}$ is also unitary, with eigenstates $\{{|\phi_{m}\rangle}^{*}\}$ and eigenvalues $\{e^{-i\phi_{m}}\}$. On the other hand, conjugating a unitary leaves its eigenvalues unchanged, so $V^{\dagger}\hat{U}V$ has the same eigenvalues $\{e^{i\phi_{m}}\}$ as $\hat{U}$. The only way a unitary $\hat{U}$ can satisfy Eq. (81), therefore, is if for every eigenvalue $e^{i\phi_{m}}$ it must also have a corresponding eigenvalue $e^{-i\phi_{m}}$. That is, every “positive energy” state must have a corresponding “negative energy” solution. (Zero-energy solutions, if any, can be taken to themselves or each other.)
We can put this condition in a more intuitive form. The inverse operator $\hat{U}^{\dagger}$ has the same eigenvectors $\{{|\phi_{m}\rangle}\}$ as $\hat{U}$, but the same eigenvalues $\{e^{-i\phi_{m}}\}$ as $\hat{U}^{*}$. We can define a new unitary operator $\tau$
@@
The existence of such a unitary $\tau$ in Eq. (83) implies the time-reversal symmetry defined in Eq. (79). Intuitively, $\tau$ is the equivalent of the classical tem reversal operation of flipping the momenta of all particles—it transforms the state so that it evolves “backwards” by transforming “positive energy” states into “negative energy” states and vice versa.
The QCAs defined in this paper all have this time-reversal symmetry, and therefore will automatically have negative-energy solutions. This creates a dilemma in how to go from free theories to interacting theories without allowing negative-energy solutions to be excited.
@@ 7 Including Interaction Terms in 1D
QCAs, as described in the earlier sections, can produce free quantum field theories in their long-wavelength (or low-energy) limit. Due to their time-reversal symmetry, these QCAs have matching negative-energy and positive-energy solutions. For free theories, the existence of negative-energy solutions has no important consequences. One can restrict the initial state to include only positive-energy states. For fermionic theories there is the additional strategy of redefining the vacuum state to be the Dirac Sea, with all negative-energy states occupied and all positive-energy states vacant. Excitation of a positive-energy state would leave a negative-energy hole: a particle-antiparticle pair. This approach has the added advantage that the dynamics can be chosen to conserve particle number.
Having two interacting QCAs, however, will raise an immediate issue. The eigenstates of $\hat{U}$—the negative- and positive-energy states—are delocalized states of definite momentum. If the interaction is local—as would be expected in a QCA—then (as we shall see) those local operators must include a non-zero component of negative-energy operators. For fermionic QCAs with a Dirac Sea vacuum this is not necessarily a problem; it merely means that interactions can give rise to particle-antiparticle pairs. But bosonic QCAs have no equivalent of the Dirac Sea; the negative energy states can have any number of excitations. So it is difficult to avoid an interaction that produces a cascade of particle-antiparticle pairs, paid for by emission of negative-energy bosons.
To simplify the analysis we will consider a 1D fermionic QCA coupled to a 1D bosonic QCA. One advantage of 1D is that it is possible to construct a fermionic QCA using only low-dimensional (qubit) local subsystems MlodinowBrun20 . The tension between locality and positive energy already exists in 1D, and the solutions are easier to find and numerically investigate.
@@ 7.1 The 1D QCAs
For our 1D QCAs, each site contains two local subsystems labeled $x,+$ and $x,-$, at positions $x=j\Delta x$ for integers $0\leq j<N$ with periodic boundaries. The evolution unitary alternately couples two qubits at neighboring sites and the two qubits at a single site, as shown in Fig. 1.
Figure 1: The structure of the 1D QCAs. For the fermionic QCA the local subsystems $x,\pm$ are qubits; for the bosonic QCA, they are harmonic oscillator modes. One step of the QCA evolution first couples the subsystems at $x,+$ and $x+\Delta x,-$ (red loops), then the subsystems at $x,+$ and $x,-$ (blue loops).
@@
In both cases, the -- sign on the ${|11\rangle}$ state produces a phase of $-1$ when two particles cross each other, producing Fermi statistics. As shown in MlodinowBrun20 one can introduce creation and annihilation operators $b_{x,\pm}^{\dagger}$ and $b_{x,\pm}$ that obey the usual anticommutation properties, and rewrite the evolution unitaries in terms of them:
$C_{x}$ $=$ $\exp\left\{-\left(\frac{\pi}{2}+\theta\right)\left(b_{x,+}^{\dagger}b_{x,-}+b_{x,-}^{\dagger}b_{x,+}\right)\right\},$
@@
We can prove that these QCAs have time-reversal symmetry by explicitly constructing the unitary operator $\tau$ in Eq. (83). Let us first make the construction in the bosonic case. Examining the unitaries defined in Eq. (90), we see that every term inside the exponential includes exactly one $+$ mode and one -- mode. If one could, for example, flip the sign of all the operators $a_{x,+}^{\dagger},a_{x,+}$, this would have the effect of flipping the sign of all the exponents, which would in turn transform $\hat{C}_{B}\rightarrow\hat{C}_{B}^{\dagger}$ and $\hat{\Sigma}_{B}\rightarrow\hat{\Sigma}_{B}^{\dagger}$. Define the unitary operator
@@ 7.2 Form of Interactions
For this paper, we will assume a particular form for this interaction; but we believe that the argument regarding locality and negative energy is very general, irrespective of the particular form of the interaction. Based on the Dirac Sea construction referred to above, we choose $U_{x}$ to conserve the number of particles in the fermionic QCA. However, based on the examples from quantum field theory, it should not conserve bosonic particle number. Using QFT as a model, we consider an interaction unitary of the form:
@@
Here, the coefficients $\{\alpha_{\ell,m,n}\}$ are parameters that represent coupling between particles in different internal states. (Generally, we will assume that these parameters are small in magnitude, $|\alpha_{\ell,m,n}|\ll 1$.) A term like the one shown in the exponential can annihilate a boson and cause a fermion to make a transition from one internal state to another; the hermitian conjugate can create a boson while causing a transition. If one thinks of states close to the Dirac Sea, where most negative-energy states are occupied and most positive-energy states are vacant, terms like this would represent a boson giving rise to a fermion-antifermion pair, or a fermion-antifermion pair annihilating to produce a boson.
However, in this local picture there is no way of knowing whether a boson produced would be positive-energy or negative-energy, and we will now see that in general the amplitude for both is nonzero.
@@ 7.3 Coupling to Negative-Energy States
The “energies” of $\hat{U}_{B}$ are ±$\pm\phi_{k}$ for $\phi_{k}=|k|\Delta x$. If we make the convenient choice of range for $k$ to be $-\pi/2\Delta x<k\leq\pi/2\Delta x$, with $k=j\pi/N\Delta x$ for some integer $-N/2<j\leq N/2$, then the result is very intuitive: the positive-energy states are $k,+$ for $k>0$ and $k,-$ for $k\leq 0$; the negative energy states are $k,-$ for $k>0$ and $k,+$ for $k\leq 0$. For each $k$ there is a positive- and a negative-energy state; we can label their annihilation operators $a_{k,\mathrm{pos}}$ and $a_{k,\mathrm{neg}}$ when we want to clarify which one we mean for a general momentum $k$.
For the fermionic QCA the “energy” eigenstates are also momentum eigenstates, but when $\theta\neq 0$ the internal basis depends on the value of $k$. However, it will suffice just to work with momentum eigenstates for the fermions.
We can use the inverse transform from Eq. (99) to rewrite the form of the interaction unitary in Eq. (98) in terms of the momentum eigenstates:
@@
where we derived this form by summing over $x$, which produces a Kronecker delta $N\delta_{k_{2},k_{1}+k_{3}}$. From the form of this interaction we can immediately conclude two things:
@@
    The coefficients $\alpha_{\ell,m,n}$ are independent of the momentum $k$; so if it is possible to create a positive-energy boson (say with $k>0$ and $\ell=+$ then it must also be possible to create a negative-energy boson (with $k<0$ and $\ell=+$).
These results are not surprising: the energy eigenstates of the boson are completely delocalized, and cannot be distinguished at a single site $x$.
@@ 7.4.1 Finite-Range Interactions
where $M$ is an even integer and $M/2$ is the range of the interaction. Since the coefficients $\alpha_{y,\ell,m,n}$ now depend on the relative position $y$, when we transform to the momentum picture the transformed coefficients can depend on momentum. Is it possible to eliminate coupling to negative-energy states?
Let’s pull out the bosonic piece of the interaction. A finite-range interaction as in Eq. (101) involves bosonic operators of the form
@@
(where we have fixed the internal states $m,n$ of the fermions and suppressed those indices). If we are allowed to choose the coefficients $\{\alpha_{y,\pm}\}$ arbitrarily, how small can we make the component of negative-energy operators in $O_{y}$ relative to the positive-energy operators? Writing the operators $a_{x+y,\pm}$ in terms of momentum operators, we get
@@
We want to choose the coefficients $\{\alpha_{y,\pm}\}$ to minimize the $\{\tilde{\alpha}_{k,\mathrm{neg}}\}$. We can quantify this by the sum
@@
where the $\{d_{jj^{\prime}}\}$ are elements of an $(M+1)\times(M+1)$ matrix $D$, and we have changed from summing over $y$ to summing over an integer index $j=1,\ldots,M+1$, where $y=(j-M/2)\Delta x$. Note that this sum is independent of $x$ (which reflects the translation symmetry of the interaction) and takes the form
@@
where $D$ is the matrix with elements $d_{jj^{\prime}}$ that we’ve just defined, and $v_{\pm}$ are vectors whose $M+1$ elements are the coefficients $\alpha_{y,\pm}$. This sum is minimized by choosing $v_{-}$ and $v_{+}$ to be the eigenvectors of $D$ and $D^{*}$, respectively, with the smallest eigenvalue. The eigenvalues of $D$ and $D^{*}$ are equal, and they must all be $\geq 0$, so we minimize the sum by choosing the eigenvalue closest to zero.
@@
The numerator takes repeating values $0,1,0,-1,0,1,\ldots$, so except for $j=j^{\prime}$ the even diagonals all vanish, and the odd diagonals take values $1/\pi$, $-1/3\pi$, $1/5\pi$, $-1/7\pi$, and so forth. Interestingly, these are the coefficients of the Taylor expansion of $(1/\pi)\arctan(1)$.
For example, for $M=4$ the matrix $\bar{D}$ is

### v2

@@ 6.1 Translation Symmetry
Translation symmetry implies that the evolution operator is invariant under shifts in the spatial lattice. In a 3D cubic lattice of size $N$ with periodic boundaries, we can define shift operators $S_{X,Y,Z}$ that move the states of the local subsystems to the site $\Delta x$ away in the positive $X$, $Y$, or $Z$ direction, respectively, wrapping around at the periodic boundaries. $S^{\dagger}_{X,Y,Z}$ obviously represents a shift in the negative direction. Translation symmetry was extensively studied in the early QCA literature Watrous95 ; Meyer96a ; Meyer96 . On the cubic lattice, these shifts all commute.
If $S$ is a product of any number of these unit shifts $S_{X,Y,Z}$ or $S^{\dagger}_{X,Y,Z}$, then the translation symmetry of the QCA evolution unitary $\hat{U}$ implies
@@
Since $\hat{U}$ and $S$ commute, we can find a basis of eigenstates of $\hat{U}$ that are eigenstates of $S$. This is what allows the momentum picture used both for QWs and QCAs in this paper.
@@ 6.3 Parity
$\mathcal{P}Q\mathcal{P}=Q,\ \ \mathcal{P}P^{\pm}_{j}\mathcal{P}=P_{j}^{\mp}.$ (78)
The simplest way of satisfying these conditions is if $\mathcal{P}$ acts on the internal space like $Q$ and if $Q$ anti-commutes with the operators $\Delta P_{j}=P_{j}^{+}-P_{j}^{-}$. This anti-commutation plays an important role in the emergence of the Dirac equation in the long-wavelength limit of QWs MlodinowBrun18 ; MlodinowBrun20 ; BrunMlodinow20 ; MlodinowBrun21 .
@@ 6.4 Time-Reversal Symmetry and Negative Energies
Time-reversal symmetry captures the idea that time evolution looks the same forward and backward. Classically, this would be illustrated by reversing the momenta of all particles while leaving their positions unchanged; this is like “running the movie backwards.” In quantum mechanics, we also need to flip the spins of particles. This symmetry is very important to the tension between locality of interactions on the one hand and positive energy on the other, as we shall see in Section 7.
@@
where $K$ is complex conjugation ($K{|\psi\rangle}={|\psi\rangle}^{*}$), and $V$ is a unitary operator. Plugging this form into Equation (79), we obtain
@@
Suppose that the evolution unitary $\hat{U}$ has eigenstates $\{{|\phi_{m}\rangle}\}$ with eigenvalues $\{e^{i\phi_{m}}\}$. Then, $\hat{U}^{*}$ is also unitary, with eigenstates $\{{|\phi_{m}\rangle}^{*}\}$ and eigenvalues $\{e^{-i\phi_{m}}\}$. On the other hand, conjugating a unitary leaves its eigenvalues unchanged, so $V^{\dagger}\hat{U}V$ has the same eigenvalues $\{e^{i\phi_{m}}\}$ as $\hat{U}$. The only way a unitary $\hat{U}$ can satisfy Equation (81), therefore, is if for every eigenvalue $e^{i\phi_{m}}$, it must also have a corresponding eigenvalue $e^{-i\phi_{m}}$. That is, every “positive-energy” solution must have a corresponding “negative-energy” solution, which leads to the existence of anti-particles. (Zero-energy solutions, if any, can be taken to themselves or each other.)
To see why $\mathcal{T}$ is interpreted as time reversal, we can put this condition in a more intuitive form. The inverse operator $\hat{U}^{\dagger}$ has the same eigenvectors $\{{|\phi_{m}\rangle}\}$ as $\hat{U}$ but the same eigenvalues $\{e^{-i\phi_{m}}\}$ as $\hat{U}^{*}$. We can define a new unitary operator $\tau$
@@
The existence of such a unitary $\tau$ in Equation (83) implies the time-reversal symmetry defined in Equation (79). Intuitively, $\tau$ is the equivalent of the classical time-reversal operation of flipping the momenta of all particles—it transforms the state so that it evolves “backwards” by transforming “positive-energy” states into “negative-energy” states and vice versa.
The QCAs defined in this paper all have this time-reversal symmetry and therefore will automatically have negative-energy solutions. We note that this conclusion holds more generally even in models that are PT-symmetric—that is, invariant under applying both a parity flip and time reversal but not under each of them separately. PT symmetry would mean that
$\hat{U}=\mathcal{PT}\hat{U}(\mathcal{PT})^{-1}.$ (84)
Since the parity-flip operator $\mathcal{P}$ is unitary, we see that $\mathcal{PT}=\mathcal{P}VK$, and we can just substitute a new unitary operator $V\rightarrow V^{\prime}\equiv\mathcal{P}V$ in Equations (8183). The argument goes through as before, and PT-symmetric models must therefore also have negative-energy solutions. This creates a dilemma in how to go from free theories to interacting theories without allowing negative-energy solutions to be excited.
@@ 7 Including Interaction Terms in 1D
QCAs, as described in the earlier sections, can produce free quantum field theories in their long-wavelength (or low-energy) limit. Due to their time-reversal symmetry, these QCAs have matching negative-energy and positive-energy solutions. For free theories, the existence of negative-energy solutions has no important consequences. One can restrict the initial state to include only positive-energy states. For fermionic theories, there is the additional strategy of redefining the vacuum state to be the Dirac Sea, with all negative-energy states occupied and all positive-energy states vacant. Excitation of a positive-energy state would leave a negative-energy hole: a particle–anti-particle pair. This approach has the added advantage that the dynamics can be chosen to conserve the particle number.
Having two interacting QCAs, however, will raise an immediate issue. The eigenstates of $\hat{U}$—the negative- and positive-energy states—are delocalized states of definite momentum. If the interaction is local—as would be expected in a QCA—then (as we shall see) those local operators must include a non-zero component of negative-energy operators. For fermionic QCAs with a Dirac Sea vacuum, this is not necessarily a problem; it merely means that interactions can give rise to particle–anti-particle pairs. But bosonic QCAs have no equivalent of the Dirac Sea; the negative-energy states can have any number of excitations. So, it is difficult to avoid an interaction that produces a cascade of particle–anti-particle pairs, paid for by the emission of negative-energy bosons.
To simplify the analysis, we will consider a 1D fermionic QCA coupled to a 1D bosonic QCA. One advantage of 1D is that it is possible to construct a fermionic QCA using only low-dimensional (qubit) local subsystems MlodinowBrun20 . The tension between locality and positive energy already exists in 1D, and the solutions are easier to find and numerically investigate.
@@ 7.1 The 1D QCAs
For our 1D QCAs, each site contains two local subsystems labeled $x,+$ and $x,-$, at positions $x=j\Delta x$ for integers $0\leq j<N$ with periodic boundaries. The evolution unitary alternately couples two qubits at neighboring sites and the two qubits at a single site, as shown in Figure 1.
Figure 1: The structure of the 1D QCAs. For the fermionic QCA, the local subsystems $x,\pm$ are qubits; for the bosonic QCA, they are harmonic oscillator modes. One step of the QCA evolution first couples the subsystems at $x,+$ and $x+\Delta x,-$ (red loops) and then the subsystems at $x,+$ and $x,-$ (blue loops).
@@
In both cases, the -- sign on the ${|11\rangle}$ state produces a phase of $-1$ when two particles cross each other, producing Fermi statistics. As shown in MlodinowBrun20 , one can introduce creation and annihilation operators $b_{x,\pm}^{\dagger}$ and $b_{x,\pm}$ that obey the usual anti-commutation properties and rewrite the evolution unitaries in terms of them:
$C_{x}$ $=$ $\exp\left\{-i\left(\frac{\pi}{2}+\theta\right)\left(b_{x,+}^{\dagger}b_{x,-}+b_{x,-}^{\dagger}b_{x,+}\right)\right\},$
@@
We can prove that these QCAs have time-reversal symmetry by explicitly constructing the unitary operator $\tau$ in Equation (83). Let us first make the construction in the bosonic case. Examining the unitaries defined in Equation (91), we see that every term inside the exponential includes exactly one $+$ mode and one -- mode. If one could, for example, flip the sign of all the operators $a_{x,+}^{\dagger},a_{x,+}$, this would have the effect of flipping the sign of all the exponents, which would in turn transform $\hat{C}_{B}\rightarrow\hat{C}_{B}^{\dagger}$ and $\hat{\Sigma}_{B}\rightarrow\hat{\Sigma}_{B}^{\dagger}$. Define the unitary operator
@@ 7.2 Form of Interactions
For this paper, we will assume a particular form for this interaction; but we believe that the argument regarding locality and negative energy is very general, irrespective of the particular form of the interaction. Based on the Dirac Sea construction referred to above, we choose $U_{x}$ to conserve the number of particles in the fermionic QCA. However, based on the examples from quantum field theory, it should not conserve the bosonic particle number. Using QFT as a model, we consider an interaction unitary of the form:
@@
Here, the coefficients $\{\alpha_{\ell,m,n}\}$ are parameters that represent coupling between particles in different internal states. (Generally, we will assume that these parameters are small in magnitude, $|\alpha_{\ell,m,n}|\ll 1$.) A term like the one shown in the exponential can annihilate a boson and cause a fermion to make a transition from one internal state to another; the hermitian conjugate can create a boson while causing a transition. If one thinks of states close to the Dirac Sea, where most negative-energy states are occupied and most positive-energy states are vacant, terms like this would represent a boson giving rise to a fermion–antifermion pair or a fermion–antifermion pair annihilating to produce a boson.
However, in this local picture, there is no way of knowing whether a boson produced would be positive-energy or negative-energy, and we will now see that, in general, the amplitude for both is non-zero.
@@ 7.3 Coupling to Negative-Energy States
The “energies” of $\hat{U}_{B}$ are ±$\pm\phi_{k}$ for $\phi_{k}=|k|\Delta x$. If we make the convenient choice of range for $k$ to be $-\pi/2\Delta x<k\leq\pi/2\Delta x$, with $k=j\pi/N\Delta x$ for some integer $-N/2<j\leq N/2$, then the result is very intuitive: the positive-energy states are $k,+$ for $k>0$ and $k,-$ for $k\leq 0$; the negative-energy states are $k,-$ for $k>0$ and $k,+$ for $k\leq 0$. For each $k$, there is a positive- and a negative-energy state; we can label their annihilation operators $a_{k,\mathrm{pos}}$ and $a_{k,\mathrm{neg}}$ when we want to clarify which one we mean for a general momentum $k$.
For the fermionic QCA, the “energy” eigenstates are also momentum eigenstates, but when $\theta\neq 0$, the internal basis depends on the value of $k$. However, it will suffice just to work with momentum eigenstates for the fermions.
We can use the inverse transform from Equation (100) to rewrite the form of the interaction unitary in Equation (99) in terms of the momentum eigenstates:
@@
where we derived this form by summing over $x$, which produces a Kronecker delta $N\delta_{k_{2},k_{1}+k_{3}}$. From the form of this interaction, we can immediately conclude two things:
@@
    The coefficients $\alpha_{\ell,m,n}$ are independent of the momentum $k$; so, if it is possible to create a positive-energy boson (say with $k>0$ and $\ell=+$), then it must also be possible to create a negative-energy boson (with $k<0$ and $\ell=+$).
These results are not surprising: the energy eigenstates of the boson are completely delocalized and cannot be distinguished at a single site $x$.
@@ 7.4.1 Finite-Range Interactions
where $M$ is an even integer and $M/2$ is the range of the interaction. Since the coefficients $\alpha_{y,\ell,m,n}$ now depend on the relative position $y$, when we transform to the momentum picture, the transformed coefficients can depend on momentum. Is it possible to eliminate coupling to negative-energy states?
Let us pull out the bosonic piece of the interaction. A finite-range interaction as in Equation (102) involves bosonic operators of the form
@@
where we have fixed the internal states $m,n$ of the fermions and suppressed those indices. If we are allowed to choose the coefficients $\{\alpha_{y,\pm}\}$ arbitrarily, how small can we make the component of negative-energy operators in $O_{y}$ relative to the positive-energy operators? Writing the operators $a_{x+y,\pm}$ in terms of momentum operators, we obtain
@@
We want to choose the coefficients $\{\alpha_{y,\pm}\}$ to minimize $\{\tilde{\alpha}_{k,\mathrm{neg}}\}$. We can quantify this by the sum
@@
where $\{d_{jj^{\prime}}\}$ are elements of an $(M+1)\times(M+1)$ matrix $D$, and we have changed from summing over $y$ to summing over an integer index $j=1,\ldots,M+1$, where $y=(j-M/2)\Delta x$. Note that this sum is independent of $x$ (which reflects the translation symmetry of the interaction) and takes the form
@@
where $D$ is the matrix with elements $d_{jj^{\prime}}$ that we have just defined, and $v_{\pm}$ are vectors whose $M+1$ elements are the coefficients $\alpha_{y,\pm}$. This sum is minimized by choosing $v_{-}$ and $v_{+}$ to be the eigenvectors of $D$ and $D^{*}$, respectively, with the smallest eigenvalue. The eigenvalues of $D$ and $D^{*}$ are equal, and they must all be $\geq 0$, so we minimize the sum by choosing the eigenvalue closest to zero.
@@
The numerator takes repeating values $0,1,0,-1,0,1,\ldots$, so except for $j=j^{\prime}$, the even diagonals all vanish, and the odd diagonals take values of $1/\pi$, $-1/3\pi$, $1/5\pi$, $-1/7\pi$, and so forth. Interestingly, these are the coefficients of the Taylor expansion of $(1/\pi)\arctan(1)$.
For example, for $M=4$, the matrix $\bar{D}$ is